OpenStudy (anonymous):

ok I have (fractions and radicals) radical 5 over 6 and I need to get it to its simplified radical form

OpenStudy (anonymous):

\[\sqrt{\frac{ 5 }{ 6 }}\]

OpenStudy (anonymous):

Simplest terms in this case means to rationalize the denominator (or in plain english - remove the radical from the denominator).

OpenStudy (anonymous):

Well I mean I need to know how! haha

OpenStudy (anonymous):

Multiplying a fraction by 1 doesn't change it's value. You can multiply the fraction by anything you want as long as the numerator equals the denominator.

OpenStudy (anonymous):

\(\sqrt6 \times \sqrt6 = 6\)

OpenStudy (anonymous):

Im so confused

OpenStudy (anonymous):

Give me about 10 minutes. First aid issue here at home.

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

Ok. back again. Poison ivy boy was scratching..... :-(

OpenStudy (anonymous):

Your fraction can be re-written as \(\dfrac{\sqrt5}{\sqrt6}\). In order to rationalize the denominator, you need to multiply the denominator by something to eliminate the square root.

OpenStudy (anonymous):

As I said earlier, when you multiply a fraction (or any term) by 1, it does not change the value. 1 can look like many different things: \(1=\dfrac{n}{n}\) or \(\dfrac{\sqrt x}{\sqrt x}\). As long as you multiply both the top and bottom of the fraction by the same thing, it will not change the value of the fraction.

OpenStudy (anonymous):

If you multiply the fraction by the square root in the denominator, you will end up with a rationalized fraction.

OpenStudy (anonymous):

That's a lot to take in, so let me know when you get back. :-)